OpenStudy (anonymous):

write an equation to solve using whole numbers and the distributive property. It costs $2.65 one way to ride the bus one way to work. He rides the bus 4 days both ways. and also buys water for $1.45 on way to work each day. How much it cost to ride and buy water. 8 bus trips and 4 waters

OpenStudy (tkhunny):

How many bus rides are there? How many water purchases were there?

OpenStudy (anonymous):

4 each way for a total of 8 trips and 4 waters

OpenStudy (tkhunny):

I see it already says that in the problem statement. Okay, write an expression for the total cost to ride the bus. Also, write an expression for the total cost of the water.

OpenStudy (anonymous):

I amsuppose to use whole numbers. 8(3.00-.35)+ 4(2.00 -.55)??? seems kinda awkward

OpenStudy (anonymous):

ex5 balloons for 2.95=5(3.00-.05)

OpenStudy (tkhunny):

How about 8*2.65 + 4*1.45 = 4(2*2.65 + 1.45)

OpenStudy (anonymous):

yes but the teacher wants it in whole numbers like my example of 5 balloons

OpenStudy (tkhunny):

Well, then do that. 8*2.65 + 4*1.45 = 8(2.00 + 0.65) + 4(1.00 + 0.45) Why is that confusing?

OpenStudy (anonymous):

it just doesn't seem any easier than 8*2.65 + 4*1.45

OpenStudy (tkhunny):

Truthfully, I'm not really seeing what is wanted. We can write it various ways, but without knowing the intent, there is little option but to guess. That's about it, then.